Find the mean of the probability distribution


of random variable X, which can take only the


values


3, 5. and 7, given that P(3)=7/30, P(5)=10/30


and P(7)=13/30.


1
Expert's answer
2022-01-12T03:33:44-0500

"M(X) = \sum {{x_i}} {p_i} = 3 \cdot \frac{7}{{30}} + 5 \cdot \frac{{10}}{{30}} + 7 \cdot \frac{{13}}{{30}} = \frac{{27}}{5}"

Answer: "\frac{{27}}{5}"